y=3(x-2)^2+3 y=6x-12
how to use substitution for these equations.

Answer by ikleyn(52778)

.
y=3(x-2)^2+3 y=6x-12
how to use substitution for these equations.
~~~~~~~~~~~~~~~~~~~~~~~~~

y = 3(x-2)^2 + 3,   (1)
y = 6x - 12.        (2)


The expression (2) for y = 6x-12 substitute into the left side of (1) insead of y. You will get

6x - 12 = 3(x_2)^2 +3.   (3)

It is single quadratic equation for x.

Simplify and solve it by any appropriate method (quadratic formula or factoring).

That's all.


You can express it in other words:

Since the left sides of equations (1) and (2) are identical, their right sides are equal:

6x - 12 = 3(x_2)^2 +3.  (4)

It is the same equation as (1).

The continuation is the same.